40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3801     Entry time: Wed Oct 27 21:51:29 2010
Author: Suresh 
Type: Update 
Category: PSL 
Subject: Attenuation of PSL NPRO removed 

The laser power was attenuated to 40 mW yesterday for ensuring that the power built up within the MC does not damage the optics. 

This however stopped us from the doubling work and besides also reduced the power available for locking the PMC. 

Therefore, today the laser attenuation was removed and once again 500mW is available at the exit of the PMC .  

However the power sent to the MC has been reduced to 60mW by changing one of the mirrors in the zig-zag to a 33% beam splitter.  Though about 450mW is incident on the beam splitter the reflected beam is only about 55mW since the mirror reflectance is specified for P polarised light incident at 45deg while ours is S-polarised incident at less than 45deg.   The light transmitted through the beam splitter has been blocked by a beam dump.

 

 

ELOG V3.1.3-